an aquarium holding african cichilds is 3in higher than it is wide.its length is 27in and its volume 9180 cubic inches. what is its height and width
Answer by Cromlix(4381)

Width = x
Height = x + 3
Length = 27 inches
x*(x + 3)* 27 = 9180 inches^3
27x(x + 3)= 9180
27x^2 + 81x - 9180 = 0
Divide throughout by 27
x^2 + 3x - 340 = 0
(x + 20)(x - 17) = 0
x = - 20 (no answer)
x = 17
So, Width = 17 inches
Height = 20 inches
Length = 27 inches.
17 * 20 * 27 = 9180
Hope this helps.
:-)